how to find no. of neutrons emitted when U undergoes .....jee 2010paper1 quest. iknow ans is 3 but is it in all d cases please tell as its not in school syllabus so....

Grade:

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

One or more neutrons within a very short time after its formation; prompt 7-rays are apparently emitted at the same time. These ideas are consistent with the results of experiments on the angular correlation between the direction of the neutrons and that of the fragments/31' the neutrons being emitted preferentially in the same direction as the fragments.

The delayed neutrons, which constitute about 0.64% of the total neutrons from the fission of U235, are emitted with gradually decreasing intensity for several minutes after the actual fission process. Six well-defined groups of delayed neutrons have been observed by studying the rate of decay of the neutron intensity/32'33' The rate of decay of each group is exponential, just as for other forms of radioactive change, and a specific half-life can be assigned to each group, as well as a mean fife and decay constant. From the intensity, the fraction 0,- which the group constitutes of the total (prompt and delayed) fission neutrons can be determined. The properties of the prominent groups of delayed neutrons are listed in Table 19-5. Three additional low-intensity delayed neutron groups from uranium have been reported/34' with half-lives of 3, 12, and 125 min and yields, per fission, of 5.8 X 10-8, 5.6 X 10~10, and 2.9 X 10-10, respectively.

vikas askiitian expert
13 years ago

235U92 => 142Xe54  +  90Sr38 + alfa particle + nuetrons       .......................1

 

at first , we find the difference of mass number of reactant to product = 235-(90+142) = 3

 

now  , we have to find difference of atomic numbers of reactant to product = 92 -(54+38) = 0

 

if difference bw atomic no is 0 as im this case , then only neutrons are emmited so reaction 1

is balanced when 3 neutrons are emmited so ans is 3 neutrons....